A constriction in a pipe reduces its diameter from 5.5 cm to 2.9 cm . Where the pipe is wider, the fluid velocity is 8 m/s. Find the fluid velocity where the pipe is narrow. Answer in units of m/s.
Answer #1

Using continuity equation,

A1 v1 = A2 v2

d1^2 v1 = d2^2 v2

0.055^2 x 8 = 0.029^2 x v2

Velocity, v2 = 28.78 m/s

Comment in case any doubt please rate my answer....